# Gradient flow for deep equilibrium single-index models

### v1

@@ 3.2.1 Gradient flow
Because $X$ is not almost surely equal to zero, we can find constants $\alpha>0$ and $\beta>0$ such that $\mathbb{E}[\lVert X\rVert_{2}^{2}\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}]\geq\beta$. Then, we may choose $\delta_{1}>\max\left\{0,\,\frac{\lVert\theta(0)\rVert_{2}}{\lVert\xi\rVert_{2}}-1\right\}$ and $\delta_{2}=\lVert\theta(0)-(\xi,0)\rVert_{2}<1/L$. Since $\sigma$ is continuously differentiable, we can find some $\gamma>0$ such that
@@ 3.2.2 Gradient descent
$\left\lVert\theta(t+1)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}$ $=\left\lVert\theta(t)-\eta\,\nabla_{\theta}\,R(\theta(t),\xi)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}$
$=\left\lVert\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}-2\eta\,\left(\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right)^{\top}\,\nabla_{\theta}\,R(\theta(t),\xi)+\eta^{2}\,\left\lVert\nabla_{\theta}\,R(\theta(t),\xi)\right\rVert_{2}^{2}.$
@@
$\begin{split}\lVert\nabla_{\theta}\,R(\theta(t),\xi)\rVert_{2}^{2}&=4\,\left\lVert\mathbb{E}\left[(y_{\theta(t)}(X)-\sigma(\xi^{\top}X))\,\frac{\sigma^{\prime}(\omega_{t}(X))}{1-\theta_{2}(t)\,\sigma^{\prime}(\omega_{t}(X))}\,\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right]\right\rVert_{2}^{2}\\ &\leq 4\,\mathbb{E}\left[(y_{\theta(t)}(X)-\sigma(\xi^{\top}X))^{2}\,\left(\frac{\sigma^{\prime}(\omega_{t}(X))}{1-\theta_{2}(t)\,\sigma^{\prime}(\omega_{t}(X))}\right)^{2}\left\lVert\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right\rVert_{2}^{2}\right].\end{split}$ (18)
@@
$(y_{\theta(t)}(X)-\sigma(\xi^{\top}X))^{2}\leq L^{2}\,((\theta_{1}(t)-\xi)^{\top}X+\theta_{2}(t)\,y_{\theta(t)}(X))^{2}\leq L^{2}\,\left\lVert\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}\,\left\lVert\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right\rVert_{2}^{2}.$
@@
$\left\lVert\theta(t+1)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}$ $\leq\left(1-\frac{2\eta\rho\gamma^{2}}{1+L\delta_{2}}+4\eta^{2}L^{2}\,\left(\frac{L}{1-L\delta_{2}}\right)^{2}\,\mathbb{E}\left[\left\lVert\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right\rVert_{2}^{4}\right]\right)\left\lVert\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}.$
@@ 4.2 Nonlinear single-index models
running gradient descent on this implicit model starting from $\theta(0)\sim\mathcal{N}(0,\,0.1^{2}\,I_{2})$ using 1000 training samples from $\mathrm{Unif}([0,1])$. For the forward pass, we use Brent’s method as the root-finding algorithm [40]. We run gradient descent for 4000 epochs with a constant step size of $\eta=0.1$. We depict the resulting loss over epochs and the learned function after training in Figure˜4, showing that gradient descent is able to successfully learn the target function.

### v2

@@ 3.2.1 Gradient flow
Because $\mathbb{E}[XX^{\top}]$ is not almost surely equal to zero, we can find a constant $\alpha>0$ such that $\mathbb{E}[XX^{\top}\,\mathbf{1}_{\lVert X\rVert_{2}\leq\alpha}]\succ 0$ by the dominated convergence theorem. Then, we may choose $\delta_{1}>\max\left\{0,\,\frac{\lVert\theta(0)\rVert_{2}}{\lVert\xi\rVert_{2}}-1\right\}$ and $\delta_{2}=\lVert\theta(0)-(\xi,0)\rVert_{2}<1/L$. Since $\sigma$ is continuously differentiable, we can find some $\gamma>0$ such that
@@ 3.2.2 Gradient descent
$\left\lVert\theta(t+1)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}$
$\quad=\left\lVert\theta(t)-\eta\,\nabla_{\theta}\,R(\theta(t),\xi)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}$
$\quad=\left\lVert\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}-2\eta\,\left(\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right)^{\top}\,\nabla_{\theta}\,R(\theta(t),\xi)+\eta^{2}\,\left\lVert\nabla_{\theta}\,R(\theta(t),\xi)\right\rVert_{2}^{2}.$
@@
$\begin{split}&\lVert\nabla_{\theta}\,R(\theta(t),\xi)\rVert_{2}^{2}\\ &\quad=4\,\left\lVert\mathbb{E}\left[(y_{\theta(t)}(X)-\sigma(\xi^{\top}X))\,\frac{\sigma^{\prime}(\omega_{t}(X))}{1-\theta_{2}(t)\,\sigma^{\prime}(\omega_{t}(X))}\,\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right]\right\rVert_{2}^{2}\\ &\quad\leq 4\,\mathbb{E}\left[(y_{\theta(t)}(X)-\sigma(\xi^{\top}X))^{2}\,\left(\frac{\sigma^{\prime}(\omega_{t}(X))}{1-\theta_{2}(t)\,\sigma^{\prime}(\omega_{t}(X))}\right)^{2}\left\lVert\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right\rVert_{2}^{2}\right].\end{split}$ (18)
@@
$(y_{\theta(t)}(X)-\sigma(\xi^{\top}X))^{2}$ $\leq L^{2}\,((\theta_{1}(t)-\xi)^{\top}X+\theta_{2}(t)\,y_{\theta(t)}(X))^{2}$
$\leq L^{2}\,\left\lVert\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}\,\left\lVert\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right\rVert_{2}^{2}.$
@@
$\left\lVert\theta(t+1)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}$
$\quad\leq\left(1-\frac{2\eta\rho\gamma^{2}}{1+L\delta_{2}}+4\eta^{2}L^{2}\,\left(\frac{L}{1-L\delta_{2}}\right)^{2}\,\mathbb{E}\left[\left\lVert\begin{pmatrix}X\\ y_{\theta(t)}(X)\end{pmatrix}\right\rVert_{2}^{4}\right]\right)\left\lVert\theta(t)-\begin{pmatrix}\xi\\ 0\end{pmatrix}\right\rVert_{2}^{2}.$
@@ 4.2 Nonlinear single-index models
running gradient descent on this implicit model starting from $\theta(0)\sim\mathcal{N}(0,\,0.1^{2}\,I_{2})$ using 1000 training samples from $\mathrm{Unif}([0,1])$. For the forward pass, we use Brent’s method as the root-finding algorithm [undefam]. We run gradient descent for 4000 epochs with a constant step size of $\eta=0.1$. We depict the resulting loss over epochs and the learned function after training in Figure˜4, showing that gradient descent is able to successfully learn the target function.